WHAT ARE "SYMMETRIC FUNCTIONS OF ROOTS"?

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Symmetric Function in Quadratic Equation

[i.e. if f(a,b) = f(b,a)].

Some of the symmetric functions of a and b are

All symmetric functions of a and b can be expressed in terms of two symmetric functions a+b and ab.

For instance,

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Aiswarya Ram Gupta
13 years ago

A function of the roots of am equation, which remains unaltered when any two of the roots are interchanged is called Symmetric function of the roots.

If a,b,c are the roots of an equation of degree 3, then we use the following notation.

∑a = a + b + c + d

∑ 1/a =1/a + 1/b + 1/c + 1/d

      ∑ ab = ab + bc + cd + da + bd + ac
∑ a2 = a2 + b+ c2 + d2

      ∑ a2b = a2b + b2a + b2c+ c2b+ c2a+ a2c+a2d+a d2 +b2 d+b d2 +c2d+ d2c

∑ a / b = a / b + b /  a + a / c + c / a + a / d + d / a + b / d + d / b + b / c + c / b + c / d +d / c